# Online POMDP Planning with Anytime Deterministic Guarantees

### v1

@@ Abstract
Autonomous agents operating in real-world scenarios frequently encounter uncertainty and make decisions based on incomplete information. Planning under uncertainty can be mathematically formalized using partially observable Markov decision processes (POMDPs). However, finding an optimal plan for POMDPs can be computationally expensive and is feasible only for small tasks. In recent years, approximate algorithms, such as tree search and sample-based methodologies, have emerged as state-of-the-art POMDP solvers for larger problems. Despite their effectiveness, these algorithms offer only probabilistic and often asymptotic guarantees toward the optimal solution due to their dependence on sampling. To address these limitations, we derive a deterministic relationship between a simplified solution that is easier to obtain and the theoretically optimal one. First, we derive bounds for selecting a subset of the observations to branch from while computing a complete belief at each posterior node. Then, since a complete belief update may be computationally demanding, we extend the bounds to support reduction of both the state and the observation spaces. We demonstrate how our guarantees can be integrated with existing state-of-the-art solvers that sample a subset of states and observations. As a result, the returned solution holds deterministic bounds relative to the optimal policy. Lastly, we substantiate our findings with supporting experimental results.
@@ 1 Introduction
Partially Observable Markov Decision Processes (POMDPs) serve as a comprehensive mathematical framework for addressing uncertain sequential decision-making problems. Despite their applicability, most problems framed as POMDPs struggle to achieve optimal solutions, largely due to factors such as large state spaces and an extensive range of potential future scenarios. The latter tends to grow exponentially with the horizon, rendering the solution process computationally prohibitive. The advent of approximate online, tree-based solvers has expanded the capacity of POMDPs, enabling them to tackle larger problems by providing a more scalable approach to problem-solving.
@@
A prominent search algorithm addressing the challenges posed by large state and observation spaces in POMDPs is POMCP (Silver and Veness, 2010). POMCP is a forward search algorithm which handles the large state and observation spaces by aggregating Monte-Carlo rollouts of future scenarios in a tree structure. During each rollout, a single state particle is recursively propagated from the root node to the leaves of the tree. It adaptively trades off between actions that lead to unexplored areas of the tree and actions that lead to rewarding areas of the tree search by utilizing UCT (Auer et al., 2002). The guarantees on the provided solution by POMCP are asymptotic, implying that the quality of the solution remains unknown within any finite time frame.
@@
While the primary focus of this paper is on discrete POMDP planning, it is essential to acknowledge recent advancements in POMDP planning that encompass both discrete and continuous observation spaces. Few notable approaches include POMCPOW (Sunberg and Kochenderfer, 2018), LABECOP (Hoerger and Kurniawati, 2021) and AdaOPS (Wu et al., 2021), which leverage explicit use of observation models. These algorithms employ importance sampling mechanisms to weigh each state sample based on its likelihood value, which is assumed to be known. Although these methods have exhibited promising performance in practical scenarios, they currently lack formal guarantees. To address this gap, (Lim et al., 2020, 2022) introduced a simplified solver aimed at bridging the theoretical gap between the empirical success of these algorithms and the absence of theoretical guarantees for continuous observation spaces. In (Lim et al., 2022), probabilistic guarantees were derived for the simplified solver concerning its proximity to the optimal value function, thus contributing to a more comprehensive understanding of POMDP planning in both discrete and continuous settings.
In this paper, we focus on deriving deterministic guarantees for POMDPs with discrete state and observation spaces. Unlike existing black-box sampling mechanisms employed in algorithms such as (Sunberg and Kochenderfer, 2018; Hoerger and Kurniawati, 2021; Wu et al., 2021), our approach assumes access not only to the observation model but also to the transition model and the prior. By leveraging this additional information, we develop a novel algorithm that utilizes a subset of the observation space, enabling the computation of deterministic bounds with respect to the optimal policy at any belief node within the constructed tree. Furthermore, we demonstrate how these deterministic bounds can be integrated with state-of-the-art algorithms, including those that sample subsets of both the state and observation spaces. In particular, we provide experimental results showcasing the use of our bounds incorporated into the AR-DESPOT algorithm.
In this paper, our main contributions are as follows. First, we derive deterministic upper and lower bounds for a POMDP problem by considering a subset of the observation space at each node along the tree under a fixed policy. Next, we extend these bounds to cover scenarios where both the state and observation spaces are restricted to subsets, enhancing the applicability of our bounds in practical settings. Based on the derived bounds, we illustrate how to incorporate the bounds into a general structure of common state-of-the-art algorithms. Specifically, in the experimental section, we present results demonstrating the integration of our theoretical guarantees with the AR-DESPOT algorithm, yielding certified solutions with deterministic bounds. Last, we perform simulations to demonstrate the use of our derived bounds in practice, further validating their relevance in POMDP planning.
@@ 2 Preliminaries
A finite horizon POMDP $M$ is defined as a tuple $\langle\mathcal{X},\mathcal{A},\mathcal{Z},T,O,\mathcal{R}\rangle$, where $\mathcal{X}$, $\mathcal{A}$, and $\mathcal{Z}$ represent a discrete state, action, and observation spaces, respectively. The transition density function $T(x_{t},a_{t},x_{t+1})\triangleq\mathbb{P}(x_{t+1}|x_{t},a_{t})$ defines the probability of transitioning from state $x_{t}\in\mathcal{X}$ to state $x_{t+1}\in\mathcal{X}$ by taking action $a_{t}\in\mathcal{A}$. The observation density function $O(x_{t},z_{t})\triangleq\mathbb{P}(z_{t}|x_{t})$ expresses the probability of receiving observation $z_{t}\in\mathcal{Z}$ from state $x_{t}\in\mathcal{X}$.
Given the limited information provided by observations, the true state of the agent is uncertain and a probability distribution function over the state space, also known as a belief, is maintained. The belief depends on the entire history of actions and observations, and is denoted $H_{t}\triangleq\{z_{1:t},a_{0:t-1}\}$. We also define the propagated history as $H_{t}^{-}\triangleq\{z_{1:t-1},a_{0:t-1}\}$. At each time step $t$, the belief is updated by applying Bayes’ rule using the transition and observation models, given the previous action $a_{t-1}$ and the current observation $z_{t}$, $b\left(x_{t}\right)=\eta_{t}\mathbb{P}(z_{t}|x_{t})\sum_{x_{t-1}\in\mathcal{X}}\mathbb{P}(x_{t}|x_{t-1},a_{t-1})b\left(x_{t-1}\right)$, where $\eta_{t}$ denotes a normalization constant and $b_{t}\triangleq\mathbb{P}(x_{t}\mid H_{t})$ denotes the belief at time t. The updated belief, $b_{t}$, is sometimes referred to as the posterior belief, or simply the posterior. We will use these interchangeably throughout the paper.
A policy function $a_{t}=\pi_{t}(b_{t})$ determines the action to be taken at time step $t$, based on the current belief $b_{t}$ and time $t$. In the rest of the paper we write $\pi_{t}\equiv\pi_{t}(b_{t})$ for conciseness. The reward is defined as an expectation over a state-dependent function, $r(b_{t},a_{t})=\mathbb{E}_{x\sim b_{t}}[r_{x}(x,a_{t})]$, and is bounded by $-\mathcal{R}_{\max}\leq r_{x}(x,a_{t})\leq\mathcal{R}_{\max}$. The value function for a policy $\pi$ over a finite horizon $\mathcal{T}$ is defined as the expected cumulative reward received by executing $\pi$ and can be computed using the Bellman update equation,
@@
The action-value function is defined by executing action $a_{t}$ and then following policy $\pi$. The optimal value function may be computed using Bellman’s principle of optimality,
@@
3 Mathematical Analysis
Typically, it is infeasible to fully expand a Partially Observable Markov Decision Process (POMDP) tree due to the extensive computational resources and time required. To address this challenge, we propose two approaches. In the first approach, we propose a solver that selectively chooses a subset of the observations to branch from, while maintaining a full posterior belief at each node. This allows us to derive a hypothetical algorithm that directly uses our suggested deterministic bounds to choose which actions to take while exploring the tree. As in some scenarios computing a complete posterior belief may be too expensive, we suggest a second method that in addition to branching only a subset of the observations, selectively chooses a subset of the states at each encountered belief. Using the deterministic bounds at the root node for each action value allows us to certify the performance of a given policy at the root of the planning tree. Moreover, given a scenario in which an action exists whose lower bound surpasses all other actions’ upper bound, the algorithm can identify the optimal action and stop further exploration. In contrast, existing state-of-the-art algorithms either do not provide any guarantee on the solution quality (e.g. POMCP Silver and Veness (2010)) or merely provide probabilistic guarantee on the solution (e.g. DESPOT Somani et al. (2013)). In the following section, we show how to use the deterministic bounds in conjunction with state-of-the-art algorithms to obtain performance guarantees.  [^1] All proofs and derivations are deferred to the supplementary file.
Both of the presented approaches diverge from many existing algorithms that rely on black-box prior, transition, and observation models. Instead, our method directly utilizes state and observation probability values to evaluate both the value function and the associated bounds. In return, we offer anytime deterministic guarantees on the value function for the derived policy and establish bounds on its deviation from the value function of the optimal policy.
In this section we provide a mathematical quantification of the impact of using a solver that only considers a small subset of the theoretical tree branches and a subset of states within each node. We begin by defining a simplified POMDP, which is a reduced complexity version of the original POMDP that abstracts or ignores certain states and observations. We then establish a connection between the simplified value function and its theoretical counterpart. Finally, we demonstrate a relationship between the simplified value function, obtained by following the best policy for the simplified problem, and the theoretically optimal value function.
We begin with general definitions of a simplified prior, transition and observation models,

### v4

# Online POMDP Planning with Anytime Deterministic Optimality Guarantees

@@ Abstract
Decision-making under uncertainty is a critical aspect of many practical autonomous systems due to incomplete information. Partially Observable Markov Decision Processes (POMDPs) offer a mathematically principled framework for formulating decision-making problems under such conditions. However, finding an optimal solution for a POMDP is generally intractable. In recent years, there has been a significant progress of scaling approximate solvers from small to moderately sized problems, using online tree search solvers. Often, such approximate solvers are limited to probabilistic or asymptotic guarantees towards the optimal solution. In this paper, we derive a deterministic relationship for discrete POMDPs between an approximated and the optimal solution. We show that at any time, we can derive bounds that relate between the existing solution and the optimal one. We show that our derivations provide an avenue for a new set of algorithms and can be attached to existing algorithms that have a certain structure to provide them with deterministic guarantees with marginal computational overhead. In return, not only do we certify the solution quality, but we demonstrate that making a decision based on the deterministic guarantee may result in superior performance compared to the original algorithm without the deterministic certification.
@@ 1 Introduction
Decision-making under uncertainty is a common challenge in many practical autonomous systems. In such systems, agents often operate with incomplete information about their environment. This uncertainty can arise from various sources, including sensor noise, hardware limitations, modeling approximations, and the inherent unpredictability of the environment. Mathematically, Decision-making under uncertainty can be formalized as Partially Observable Markov Decision Process (POMDP).
Unfortunately, finding an optimal solution to most POMDP problems is computationally intractable, mostly due to a large number of possibilities for the ground truth of the current state, and exponentially increasing possibilities of the future outcomes, commonly referred to as the curse of dimensionality, and the curse of history. As such, most state-of-the-art (SOTA) algorithms aim to find an approximate solution.
One prominent approach to deriving approximate solutions employs an online tree-search paradigm. In this framework, following each real-world decision, an online solver evaluates the current state and projects potential future scenarios. These scenarios are organized within a tree graph structure. As the tree is constructed, the agent assesses the implications of selecting a particular action, subsequently receiving feedback from the environment. This feedback informs the estimation of probabilities for new states, guiding the selection of subsequent actions based on accumulated knowledge. This iterative process continues, building on past outcomes to navigate the decision space.
Given the inherent approximation in these solutions, a natural inquiry regarding the connection between the approximate solution and the actual problem at hand. Some state of the art online algorithms, e.g. Silver and Veness (2010), offer asymptotic guarantees thus having no finite time guarantees on the solution quality. A different class of algorithms suggests finite time, but probabilistic guarantees such as Somani et al. (2013). Many algorithms have shown good empirical performance, at the advent of the practical use case of POMDP problems, e.g. Sunberg and Kochenderfer (2018), but fall short of providing a framework that bridges between the policy found and the underlying POMDP.
In this paper, we focus on deriving deterministic guarantees for POMDPs with discrete state, action and observation spaces. Unlike existing black-box sampling mechanisms employed in algorithms such as (Sunberg and Kochenderfer, 2018; Hoerger and Kurniawati, 2021; Wu et al., 2021), our approach assumes access not only to the observation model but also to the transition and the prior models. By leveraging this additional information, we develop novel bounds that necessitate only a subset of the state and observation spaces, enabling the computation of deterministic bounds with respect to the optimal policy at any belief node within the constructed tree. From a practical standpoint, we demonstrate how to harness the theoretical derivations to recent advancements in POMDP approximate solvers, by attaching the bounds to existing state-of-the-art algorithms. We show that despite their stochastic nature, we can guarantee deterministic linkage to the optimal solution with marginal computational overhead. We extend the approach even further by demonstrating how to utilize the bounds to explore the tree and finally select an action based on the deterministic guarantees.
In this paper, our main contributions are as follows. First, we introduce a simplified POMDP that uses a subset of the state and observation spaces to increase the computational efficiency. Then, we derive deterministic bounds that relate between the former and the non-simplified POMDP. Notably, the bounds are only a function of the states and observations known to the simplified POMDP and hence can be calculated while planning to guide the decision-making and even exploration. We also show a tighter version of the bounds considered in the conference version of this paper, (Barenboim and Indelman, 2023). We further extend the approach and show that utilizing these bounds for exploration results in convergence to the optimal solution of the POMDP in finite time; While the optimality guarantees applied only to observation-space simplification in the conference version, we extend the results in this paper by deriving optimality guarantees for both state- and observation-space simplification. Based on the derived bounds, we illustrate how to incorporate the bounds into a general structure of common state-of-the-art algorithms. We utilize the bounds for exploration, decision-making and pruning of suboptimal actions while planning. Last, we demonstrate the practicality of the bounds by experimenting with our novel algorithms, suggested in this paper, namely DB-POMCP, RB-POMCP and DB-DESPOT, which are variants of the POMCP and DESPOT algorithms, to improve the empirical results in finite-horizon problems.
2 Related Work
Over the last two decades there has been significant progress in online POMDP planning, aiming to balance the trade-off between computational efficiency and the quality of the solution.
The Heuristic Search Value Iteration (HSVI) Smith and Simmons (2004) algorithm marked a significant milestone in POMDP planning by introducing an efficient point-based value iteration method that provides convergence guarantees. HSVI leverages a heuristic to focus the search on the most promising regions of the belief space, thus improving computational efficiency while maintaining solution quality. Another pivotal algorithm, Successive Approximations of the Reachable Space under Optimal Policies (SARSOP) Kurniawati et al. (2008), builds on this idea by further refining the focus on reachable belief states under optimal policies. SARSOP’s ability to prune irrelevant parts of the belief space enables it to handle larger POMDPs more effectively. However, these approaches were limited in their scalability to large state spaces due to the necessity of computing a complete belief state at each posterior node in the planning tree.
@@ 2 Related Work
The advent of Monte Carlo methods brought a significant shift in online POMDP planning. The Partially Observable Monte Carlo Planning (POMCP) (Silver and Veness, 2010) algorithm introduced a particle filter-based approach combined with Monte Carlo tree search (MCTS). POMCP uses a set of particles to represent the belief state and UCT (Upper Confidence bounds applied to Trees, (Couëtoux et al., 2011)) to guide the search, making it much more scalable and for large state and observation spaces. POMCP is a forward search algorithm which handles the large state and observation spaces by aggregating Monte-Carlo rollouts of future scenarios in a tree structure. During each rollout, a single state particle is recursively propagated from the root node to the leaves of the tree. It adaptively trades off between actions that lead to unexplored areas of the tree and actions that lead to rewarding areas of the tree search by utilizing UCT (Auer et al., 2002). The guarantees on the provided solution by POMCP are asymptotic, implying that the quality of the solution remains unknown within any finite time frame.
@@
While the primary focus of this paper is on discrete POMDP planning, it is essential to acknowledge recent advancements in POMDP planning that encompass both discrete and continuous observation spaces. Few notable approaches include POMCPOW (Sunberg and Kochenderfer, 2018), LABECOP (Hoerger and Kurniawati, 2021) and AdaOPS (Wu et al., 2021), which leverage explicit use of observation models. These algorithms employ importance sampling mechanisms to weigh each state sample based on its likelihood value, which is assumed to be known. Although these methods have exhibited promising performance in practical scenarios, they currently lack formal guarantees. To address this gap, (Lim et al., 2020, 2022) introduced a simplified solver aimed at bridging the theoretical gap between the empirical success of these algorithms and the absence of theoretical guarantees for continuous observation spaces. In (Lim et al., 2022), probabilistic guarantees were derived for the simplified solver estimating its proximity to the optimal value function.
@@ 3 Preliminaries
A finite horizon POMDP $M$ is defined as a tuple $\langle\mathcal{X},\mathcal{A},\mathcal{Z},T,O,\mathcal{R},b_{0}\rangle$, where $\mathcal{X}$, $\mathcal{A}$, and $\mathcal{Z}$ represent a discrete state, action, and observation spaces, respectively. The transition density function $T(x_{t},a_{t},x_{t+1})\triangleq\mathbb{P}(x_{t+1}|x_{t},a_{t})$ defines the probability of transitioning from state $x_{t}\in\mathcal{X}$ to state $x_{t+1}\in\mathcal{X}$ by taking action $a_{t}\in\mathcal{A}$. The observation density function $O(x_{t},z_{t})\triangleq\mathbb{P}(z_{t}|x_{t})$ expresses the probability of receiving observation $z_{t}\in\mathcal{Z}$ from state $x_{t}\in\mathcal{X}$. $b_{0}\equiv\mathbb{P}(x_{0}\mid H_{0})$ represents the prior probability function, which is the distribution function over the state space at time $t=0$.
Given the limited information provided by observations, the true state of the agent is uncertain and a probability distribution function over the state space, also known as a belief, is maintained. The belief depends on the entire history of actions and observations, denoted as $H_{t}\triangleq\{z_{1:t},a_{0:t-1}\}$. We also define the propagated history as $H_{t}^{-}\triangleq\{z_{1:t-1},a_{0:t-1}\}$. At each time step $t$, the belief is updated by applying Bayes’ rule using the transition and observation models, given the previous action $a_{t-1}$ and the current observation $z_{t}$, $b\left(x_{t}\right)=\eta_{t}\mathbb{P}(z_{t}|x_{t})\sum_{x_{t-1}\in\mathcal{X}}\mathbb{P}(x_{t}|x_{t-1},a_{t-1})b\left(x_{t-1}\right)$, where $\eta_{t}$ denotes a normalization constant and $b_{t}\triangleq\mathbb{P}(x_{t}\mid H_{t})$ denotes the belief at time t. The updated belief, $b_{t}$, is sometimes referred to as the posterior belief, or simply the posterior. We will use these interchangeably throughout the paper.
A policy function $a_{t}=\pi_{t}(H_{t})$ determines the action to be taken at time step $t$, based on the history $H_{t}$ and time $t$. In the rest of the paper we write $\pi_{t}\equiv\pi_{t}(H_{t})$ for conciseness. The reward is defined as an expectation over a state-dependent function, $r(b_{t},a_{t})=\mathbb{E}_{x\sim b_{t}}[r_{x}(x,a_{t})]$, and is assumed to be bounded by $-\mathcal{R}_{\max}\leq r_{x}(x,a_{t})\leq\mathcal{R}_{\max}$. The value function for a policy $\pi$ over a finite horizon $T$ is defined as the expected cumulative reward received by executing $\pi$ and can be computed using the Bellman update equation,
@@
We use $V^{\pi}_{t}(b_{t})$ and $V^{\pi}_{t}(H_{t})$ interchangeably throughout the paper. The action-value function is defined by executing action $a_{t}$ and then following policy $\pi$,
$Q^{\pi}_{t}(b_{t},a_{t})=r(b_{t},a_{t})+\underset{z_{t+1:T}}{\mathbb{E}}\left[\sum_{\tau=t+1}^{T}r(b_{\tau},\pi_{\tau})\right].$ (2)
The optimal value function may be computed using Bellman’s principle of optimality,
@@
For notational convenience, we introduce a few more simplifying notations; We use $\mathcal{V}_{max,t},\mathcal{V}_{min,t}$ to denote upper an lower bounds on the value function at time step $t$. In the simplest case, these may be $\mathcal{V}_{max,t}=(T-t)\cdot\mathcal{R}_{\max}$, $\mathcal{V}_{min,t}=(t-T)\cdot\mathcal{R}_{\max}$. Additionally, we denote a trajectory as, $\tau_{t}=\{x_{0},a_{0},z_{1},x_{1},a_{1},\dots,a_{t-1},x_{t},z_{t}\}$, and a corresponding probability distribution over the possible trajectories, $\mathbb{P}(\tau_{t})$. We denote a policy-dependent trajectory distribution as $\mathbb{P}^{\pi}(\tau_{t})\equiv\mathbb{P}(\tau_{t}\mid b_{0},\pi_{0},\dots,\pi_{t})$.
4 Simplified POMDP
Typically, it is infeasible to fully expand a Partially Observable Markov Decision Process (POMDP) tree due to the extensive computational resources and time required. To address this challenge, we propose two approaches. In the first approach, presented in 5.1, we propose a solver that selectively chooses a subset of the observations to branch from, while maintaining a full posterior belief at each node. This allows us to derive an hypothetical algorithm that directly uses our suggested deterministic bounds to choose which actions to take while exploring the tree. As in most scenarios computing a complete posterior belief may be too expensive, in section 5.2 we suggest an improved method that in addition to branching only a subset of the observations, selectively chooses a subset of the states at each encountered belief.
The presented approaches diverge from many existing algorithms that rely on black-box prior, transition, and observation models. Instead, our method directly utilizes state and observation probability values to evaluate both the value function and the associated bounds. In return, an anytime deterministic guarantee on the value function for the derived policy concerning its deviation from the optimal value function is derived.
To that end, we define a simplified POMDP, which is a reduced version of the original POMDP that abstracts or ignores certain states and/or observations. A simplified POMDP, $\bar{M}$, is a tuple $\langle\bar{\mathcal{X}},\mathcal{A},\bar{\mathcal{Z}},\bar{T},\bar{O},\mathcal{R},\bar{b}_{0}\rangle$, where $\bar{\mathcal{X}},\bar{\mathcal{Z}},\bar{T}$ and $\bar{O}$ are the simplified versions of the state and observation spaces, and their corresponding transition and observation models,